OpenStudy (anonymous):

Factor the expression given below. Write each factor as a polynomial in descending order. Enter exponents using the caret ( ^ ). For example, you would enter x2 as x^2. 125x3 + 216y3

OpenStudy (magbak):

@haiimawesome

OpenStudy (anonymous):

@magbak HELP :D

OpenStudy (dan815):

i think uhh

OpenStudy (magbak):

I have a solver site you want it it calculates it for you

OpenStudy (magbak):

http://www.mathportal.org/calculators.php

OpenStudy (magbak):

125x^{3}+216y^{3} = (5x+6y)(25x^{2}-30xy+36y^{2}) that's the answer

OpenStudy (dan815):

\[a^3-b^3=(a-b)(a^2+ab+b^2)\] \[5^3a^3-((-6)^3b^3)=(5a-(-6b)(25a^2+5*-6ab+36*b^2)\]

OpenStudy (anonymous):

THANK YOU GUYS

OpenStudy (dan815):

and it simplifies to magbaks solution

OpenStudy (dan815):

@haiimawesome do you understand how this happened?

OpenStudy (anonymous):

yes, thank you @dan815

OpenStudy (anonymous):

@magbak

OpenStudy (magbak):

Its B

OpenStudy (anonymous):

mkay haha thankkkkkkssss <3

OpenStudy (magbak):

<3

OpenStudy (anonymous):

@magbak

OpenStudy (anonymous):

@chucho78

OpenStudy (anonymous):

hmm?

OpenStudy (anonymous):

the attatched screenshop

OpenStudy (anonymous):

*screenshot

OpenStudy (anonymous):

this one

OpenStudy (anonymous):

@Lets_Be_Cops @mkerr017